# Agentic Multi-Persona Framework for Evidence-Aware Fake News Detection

### v1

@@ 4.4. LLM Pseudo-Labeling: Policy, Decision Rules, and Output
(6) $\scriptsize(\mathcal{Y}_{p},\,\mathcal{C}_{s},\,\mathcal{J})\;=\;\Phi~\!\bigl(~\pi_{c};\;\mathcal{M}^{\tau},\,{\text{FewShots}}~\bigr),$
@@
4.4.1. Persuasion-based Refinement
@@
(7) $\scriptsize\mathcal{U}=\{N\mid\mathcal{Y}_{p}(N)=\textit{UNCERTAIN}=\Phi~\!\bigl(~\pi_{c};\;\mathcal{M}^{\tau},\,\text{FewShots}~\bigr)\}.$
@@
Action space: $\mathcal{A}_{\text{PKM}}=\Bigl\{\,z\in\{0,1\}^{T},~\ \mathcal{S}=\{s_{t}\},~\ u\in\mathbb{N}^{G},~\ \Pi\in[0,1],~\ \mathcal{J}_{\text{PKM}}\Bigr\}$. Here, $T=23$ persuasion techniques (JRC taxonomy (piskorski2023news)) and $G=6$ categories (justification, simplification, distraction, call, reputation attack, and manipulative wording); $z$ is the technique-activation vector, $\mathcal{S}$ is extracted evidence spans (phrases), $u$ is category counts, $\Pi$ is a normalized persuasion index, and $\mathcal{J}_{\text{PKM}}$ is a concise narrative summary.
Policy: $\pi_{\text{PKM}}:\ \mathcal{O}\times\mathcal{M}\ \longrightarrow\ \mathcal{A}_{\text{PKM}}$. Given $(\mathcal{C},\mathcal{T}_{p},\mathcal{I}_{s},\mathcal{M})$, the policy produces (i) $z,\mathcal{S}$ via detection over content; (ii) $u=\mathscr{A}z$ using a fixed category–technique incidence matrix ${\mathscr{A}}\in\{0,1\}^{G\times T}$; (iii) $\Pi=\frac{\beta^{\top}u}{\|\beta\|_{1}}$ with $\beta\in\mathbb{R}_{+}^{G}$ denoting the importance of the category; (iv) $\mathcal{J}_{\text{PKM}}=\mathrm{Summarize}(z,\mathcal{S})$. The PKM agent is gated, it is invoked only if ${g_{\text{PKM}}}={\mathbb{1}}[\mathcal{Y}_{p}=\textit{UNCERTAIN}]$; $\mathbb{1}$}[.] is the indicator function.
@@
4.4.2. LLM Pseudo-Labeling Policy
@@
${\Phi}:\ \mathcal{M}^{\tau}\ \mapsto\ ({\mathcal{Y}}_{p},{\mathcal{C}}_{s},{\mathcal{J}})~~;\quad\tilde{\Phi}:\ \mathcal{M}^{+}\mapsto\ (\mathcal{Y}_{p},\mathcal{C}_{s},\mathcal{J}),\text{ if }g_{\text{PKM}}=1$
@@
(8) $\scriptsize S^{+}=\sum_{d_{i}\in\mathcal{E}^{+}}\mathcal{R}(d_{i});\qquad S^{-}=\sum_{d_{i}\in\mathcal{E}^{-}}\mathcal{R}(d_{i});\qquad$
@@
(9) $\scriptsize\tilde{\mathcal{Y}}_{p}\;=\;\begin{cases}\textit{REAL},&\left(|\mathcal{E}^{+}_{H}|>0\wedge|\mathcal{E}^{-}_{H}|=0\right)\ \ \text{or}\\[1.0pt] &\left(|\mathcal{E}^{+}_{H}|+|\mathcal{E}^{-}_{H}|=0\wedge S^{+}_{\text{Med}}-S^{-}_{\text{Med}}\geq\gamma\right),\\[4.0pt] \textit{FAKE},&\left(|\mathcal{E}^{-}_{H}|>0\wedge|\mathcal{E}^{+}_{H}|=0\right)\ \ \text{or}\\[1.0pt] &\left(\text{only low-reliable support}\right)\ \ \text{or}\ \ \left(\psi_{\text{impl}}(N)=1\right),\\[4.0pt] \textit{UNCERTAIN},&\left(|\mathcal{E}^{+}_{H}|>0\wedge|\mathcal{E}^{-}_{H}|>0\right)\ \ \text{or}\ \ \left(S^{+}+S^{-}<\eta\right),\end{cases}$
@@
4.5. SLM Classifier for Final Output
@@
(10) $\scriptsize x=\big[\mathcal{H}\;\oplus\;\langle\texttt{SEP}\rangle\;\mathcal{T}_{p}\;\oplus\;\langle\texttt{SEP}\rangle\;\mathcal{I}_{s}\;\oplus\;\langle\texttt{SEP}\rangle\;\mathcal{J}\big],$
@@ 5. Experiments
All experiments were conducted on a system equipped with PyTorch 2.5.1, CUDA 12.1, NVIDIA A100-SXM4-40GB GPU, AMD EPYC 7742 64-core CPU with 1 TB RAM. We used the Google Cloud Custom Search API to retrieve external evidence, and the OpenAI GPT-4o mini for LLM-based inference. For the final verdict, DistilRoBERTa-base (distilroberta-base) was finetuned as SLM to obtain REAL or FAKE.
5.1. Datasets and Evaluation Metrics
@@
Datasets: To evaluate our model, we used three widely adopted fake news datasets: PolitiFact (shu2020fakenewsnet), GossipCop (shu2020fakenewsnet), and MMCoVaR (chen2021mmcovar). PolitiFact contains fact-checked political claims with veracity labels, while GossipCop consists of celebrity and entertainment-related news articles, annotated as REAL or FAKE. MMCoVaR is a multimodal dataset on COVID-19 vaccine-related misinformation, comprising both textual posts and associated images. The detailed statistics of all datasets are presented in Table 1. Brief statistics of these datasets are presented in Table 1. For training and evaluation, we used the train-validation-test splits in a ratio of 7:2:17{:}2{:}1.
Table 1. Statistics of the datasets
Statistics PolitiFact GossipCop MMCoVaR
# Fake News 103 250 706
# Real News 172 250 1530
# Images 275 500 2236
@@
Evaluation Metrics: To evaluate the effectiveness of the model, we employ four widely adopted metrics: Accuracy % ($Acc$), F1 Score % ($F1$), Precision % ($Pre$), and Recall % ($Rec$). In addition, we report an improvement metric that measures the percentage gain in our method over the second-best baseline. To evaluate the quality of the generated rationale (LLM justification $\mathcal{J}$), we employ three complementary metrics: (i) the Utility metric $\mathcal{U}$, capturing the effectiveness of the rationale as the performance difference between fine-tuned SLMs trained with and without LLM justification $\mathcal{J}$; (ii) the Leakage-Adjusted Simulatability ($LAS$) metric (las_metric), assessing faithfulness, how well $\mathcal{J}$ aids in simulating the verdict without leaking label information, computed as $\mathbb{1}[\mathcal{Y}_{s}\mid\mathcal{H},\mathcal{T}_{p},\mathcal{I}_{s},\mathcal{J}]-\mathbb{1}[\mathcal{Y}_{s}\mid\mathcal{H},\mathcal{T}_{p},\mathcal{I}_{s}];$ where $\mathbb{1}[\cdot]$ is an indicator function that equals $1$ if the model correctly predicts the true label and 0 otherwise; and (iii) the Rationale Evaluation via conditional $\mathcal{V}$-information ($REV$) metric (rev_metric), quantifying informativeness, the additional label-relevant information provided by $\mathcal{J}$, computed as the log-likelihood difference of the true label $\mathcal{Y}$ between two models $f_{\theta}$ and $f^{\prime}_{\theta}$ trained with $\mathcal{J}$ and without $\mathcal{J}$ as inputs, respectively: $\log f_{\theta}([\mathcal{H},\mathcal{T}_{p},\mathcal{I}_{s}],\mathcal{J})(\mathcal{Y})-\log f^{\prime}_{\theta}([\mathcal{H},\mathcal{T}_{p},\mathcal{I}_{s}])(\mathcal{Y}).$ Each metric is averaged over all samples in the test dataset.
5.2. Results and Analysis
Table 2. Comparison with SOTA methods
Dataset Method AccAcc $F1$ PrePre RecRec PolitiFact LMDD (Kheddache2025) 80.86 75.99 80.05 72.34 FactAgent (Li2024) 81.41 81.09 80.79 81.41 MFCL (CHEN2025112800) 81.82 82.52 81.61 83.47 BREAK (yin2025break) 82.14 82.03 81.79 82.29 MGCA (Guo_Ma_Zeng_Luo_Zeng_Tang_Zhao_2025) 82.14 81.10 82.14 80.09 EARAM (10.1145/3696410.3714532) 82.93 81.11 80.90 81.32 FakingRecipe (10.1145/3664647.3680663) 83.33 80.07 80.83 79.44 AMPEND-LS (Proposed) 92.18 ± 1.93 91.64 ± 2.12 92.06 ± 1.85 91.40 ± 2.43 Improvement (%) 10.62 11.05 12.08 9.50 GossipCop LMDD (Kheddache2025) 65.17 71.05 78.21 65.10 FactAgent (Li2024) 78.04 77.95 78.45 77.50 FakingRecipe (10.1145/3664647.3680663) 80.26 80.27 80.21 80.26 MFCL (CHEN2025112800) 81.81 81.35 81.65 81.06 BREAK (yin2025break) 82.00 82.11 82.51 81.73 EARAM (10.1145/3696410.3714532) 85.00 85.49 86.00 85.00 MGCA (Guo_Ma_Zeng_Luo_Zeng_Tang_Zhao_2025) 88.00 84.09 88.31 80.26 AMPEND-LS (Proposed) 93.43 ± 1.53 93.24 ± 1.59 93.44 ± 1.48 93.29 ± 1.55 Improvement (%) 6.17 9.07 5.81 9.75 MMCoVaR LMDD (Kheddache2025) 64.62 63.54 78.61 53.40 FactAgent (Li2024) 79.14 73.70 74.45 73.10 FakingRecipe (10.1145/3664647.3680663) 79.96 78.61 79.35 77.92 MFCL (CHEN2025112800) 82.06 78.19 77.82 78.57 MGCA (Guo_Ma_Zeng_Luo_Zeng_Tang_Zhao_2025) 90.00 88.34 88.19 88.48 EARAM (10.1145/3696410.3714532) 90.36 88.51 88.32 88.70 BREAK (yin2025break) 91.89 89.27 88.73 89.83 AMPEND-LS (Proposed) 96.46 ± 1.11 95.56 ± 1.31 96.30 ± 1.48 94.96 ± 1.55 Improvement (%) 4.97 7.05 8.53 5.71
@@ 5.2. Results and Analysis
To benchmark our approach, AMPEND-LS, we compare it against two sets of baselines, (i) LLM-based methods: LMDD (Kheddache2025) and FactAgent (Li2024), and (ii) Learning-based methods, including five recent state-of-the-art (SOTA) methods, FakingRecipe (10.1145/3664647.3680663), MFCL (CHEN2025112800), MGCA (Guo_Ma_Zeng_Luo_Zeng_Tang_Zhao_2025), BREAK (yin2025break), and EARAM (10.1145/3696410.3714532).
@@
Table 2 shows that AMPEND-LS consistently outperformed the SOTA baselines in all three datasets: PolitiFact, GossipCop, and MMCoVaR, achieving the highest accuracy, $F1$, precision, and recall simultaneously. The largest improvements were observed on PolitiFact (+10.62% $Acc$, +11.84% $F1$), highlighting its ability to handle politically nuanced misinformation, while steady improvements on GossipCop and MMCoVaR confirm strong generalization to both text-heavy and multimodal inputs. Unlike baselines that trade off precision (e.g., MGCA) or recall (e.g., MFCL), AMPEND-LS achieved balanced improvements across all metrics, attributable to its complementary modules of agentic reasoning, persuasion-based refinement, and credibility-guided evidence fusion. Competing methods, such as BREAK, performed well only in specific domains (e.g., MMCoVaR) but failed to generalize. The tight alignment of accuracy, precision, recall, and $F1$ in AMPEND-LS further indicates stable decision boundaries and reduced overfitting. Collectively, these results validate AMPEND-LS as a robust, adaptable, and evidence-grounded framework capable of handling diverse misinformation domains with transparency and resilience.
5.3. Ablation Study
To assess the contribution of individual components in AMPEND-LS, we conduct a systematic ablation on PolitiFact and GossipCop, with results shown in Table 3. The key findings are:
Table 3. Module-wise ablation study

### v2

@@ IV-D LLM Pseudo-Labeling: Policy, Decision Rules, and Output
$\small(\mathcal{Y}_{p},\mathcal{C}_{s},\mathcal{J})=\Phi~\bigl(\pi_{c};\mathcal{M}^{\tau},{\text{FewShots}}\bigr)$ (6)
@@
IV-D1 Persuasion-based Refinement
@@
$\scriptsize\mathcal{U}=\{N\mid\mathcal{Y}_{p}(N)=\textit{UNCERTAIN}=\Phi~\bigl(~\pi_{c};\mathcal{M}^{\tau},\text{FewShots}~\bigr)\}.$ (7)
@@
Action space: $\mathcal{A}_{\text{PKM}}=\Bigl\{z\in\{0,1\}^{T},~\ \mathcal{S}=\{s_{t}\},~\ u\in\mathds{N}^{G},~\ \Pi\in[0,1],~\ \mathcal{J}_{\text{PKM}}\Bigr\}$. Here, $T=23$ persuasion techniques (JRC taxonomy [19]) and $G=6$ categories (justification, simplification, distraction, call, reputation attack, and manipulative wording); $z$ is the technique-activation vector, $\mathcal{S}$ is extracted evidence spans (phrases), $u$ is category counts, $\Pi$ is a normalized persuasion index, and $\mathcal{J}_{\text{PKM}}$ is a concise narrative summary.
[…]
@@
IV-D2 LLM Pseudo-Labeling Policy
@@
${\Phi}:\ \mathcal{M}^{\tau}\ \mapsto\ ({\mathcal{Y}}_{p},{\mathcal{C}}_{s},{\mathcal{J}})~;~~\tilde{\Phi}:\ \mathcal{M}^{+}\mapsto\ (\mathcal{Y}_{p},\mathcal{C}_{s},\mathcal{J}),\text{ if }g_{\text{PKM}}=1$
@@
$\scriptsize S^{+}=\sum_{d_{i}\in\mathcal{E}^{+}}\mathcal{R}(d_{i})~;~~S^{-}=\sum_{d_{i}\in\mathcal{E}^{-}}\mathcal{R}(d_{i}),$ (8)
@@
$\scriptsize\tilde{\mathcal{Y}}_{p}=\begin{cases}\textit{REAL},&\left(|\mathcal{E}^{+}_{H}|>0\wedge|\mathcal{E}^{-}_{H}|=0\right)\ \ \text{or}\\[1.0pt] &\left(|\mathcal{E}^{+}_{H}|+|\mathcal{E}^{-}_{H}|=0\wedge S^{+}_{\text{Med}}-S^{-}_{\text{Med}}\geq\gamma\right);\\[4.0pt] \textit{FAKE},&\left(|\mathcal{E}^{-}_{H}|>0\wedge|\mathcal{E}^{+}_{H}|=0\right)\ \ \text{or}\\[1.0pt] &\left(\text{only low-reliable support}\right)\ \ \text{or}\ \ \left(\psi_{\text{impl}}(N)=1\right);\\[4.0pt] \textit{UNCERTAIN},&\left(|\mathcal{E}^{+}_{H}|>0\wedge|\mathcal{E}^{-}_{H}|>0\right)\ \ \text{or}\ \ \left(S^{+}+S^{-}<\eta\right)\end{cases}$ (9)
@@
IV-E SLM Classifier for Final Output
@@
$\scriptsize x=\big[\mathcal{H}\oplus\langle\texttt{SEP}\rangle\mathcal{T}_{p}\oplus\langle\texttt{SEP}\rangle\mathcal{I}_{s}\oplus\langle\texttt{SEP}\rangle\mathcal{J}\big],$ (10)
@@ V Experiments
All experiments were conducted on a system equipped with PyTorch 2.5.1, CUDA 12.1, NVIDIA A100-SXM4-40GB GPU, AMD EPYC 7742 64-core CPU with 1 TB RAM. We used the Google Cloud Custom Search API to retrieve external evidence, and the OpenAI GPT-4o mini for LLM-based inference. For the final verdict, DistilRoBERTa-base [22] was finetuned as SLM to obtain REAL or FAKE.
V-A Datasets and Evaluation Metrics
V-A1 Datasets
@@
To evaluate our model, we used three widely adopted fake news datasets: PolitiFact [25], GossipCop [25], and MMCoVaR [4]. PolitiFact contains fact-checked political claims with veracity labels, while GossipCop consists of celebrity and entertainment news articles annotated as REAL or FAKE. MMCoVaR is a multimodal dataset focused on COVID-19 vaccine-related misinformation, comprising textual posts along with associated images. We adopted a 7:2:17{:}2{:}1 train–validation–test split, and all reported results are based on the test set.
V-A2 Evaluation Metrics
@@
To evaluate the effectiveness of the model, we employ four widely adopted metrics: Accuracy % ($Acc$), F1 Score % ($F1$), Precision % ($Pre$), and Recall % ($Rec$). In addition, we report an improvement metric that measures the percentage gain in our method over the second-best baseline. To evaluate the quality of the generated rationale (LLM justification $\mathcal{J}$), we employ three complementary metrics: (i) the Utility metric $\mathcal{U}$, capturing the effectiveness of the rationale as the performance difference between fine-tuned SLMs trained with and without LLM justification $\mathcal{J}$; (ii) the Leakage-Adjusted Simulatability ($LAS$) metric [10], assessing faithfulness, how well $\mathcal{J}$ aids in simulating the verdict without leaking label information, computed as $\mathds{1}[\mathcal{Y}_{s}\mid\mathcal{H},\mathcal{T}_{p},\mathcal{I}_{s},\mathcal{J}]-\mathds{1}[\mathcal{Y}_{s}\mid\mathcal{H},\mathcal{T}_{p},\mathcal{I}_{s}]$, where $\mathds{1}[\cdot]$ is an indicator function that equals $1$ if the model correctly predicts the true label and 0 otherwise; and (iii) the Rationale Evaluation via conditional $\mathcal{V}$-information ($REV$) metric [3], quantifying informativeness, the additional label-relevant information provided by $\mathcal{J}$, computed as the log-likelihood difference of the true label $\mathcal{Y}$ between two models $f_{\theta}$ and $f^{\prime}_{\theta}$ trained with $\mathcal{J}$ and without $\mathcal{J}$ as inputs, respectively: $\log f_{\theta}([\mathcal{H},\mathcal{T}_{p},\mathcal{I}_{s}],\mathcal{J})(\mathcal{Y})-\log f^{\prime}_{\theta}([\mathcal{H},\mathcal{T}_{p},\mathcal{I}_{s}])(\mathcal{Y})$.
V-B Results and Analysis
@@
To benchmark our approach, AMPEND-LS, we compare it against two sets of baselines, (i) LLM-based methods: LMDD [13] and FactAgent [16], and (ii) Learning-based methods, including five recent state-of-the-art (SOTA) methods, MFCL [2], BREAK [36], MGCA [8], EARAM [40], and FakingRecipe [1].
@@
As shown in Table I, AMPEND-LS consistently outperforms all baselines across the three datasets—PolitiFact, GossipCop, and MMCoVaR—achieving the highest accuracy, $F1$, precision, and recall in each case. The most substantial gains are observed on GossipCop, with improvements of +6.17% in accuracy and +9.07% in $F1$ over the strongest baseline, demonstrating the framework’s effectiveness in handling entertainment-oriented and rumor-driven content. On PolitiFact and MMCoVaR, steady improvements further confirm strong generalization across both text-centric and multimodal misinformation settings. Unlike several competing methods that exhibit imbalanced precision–recall trade-offs (e.g., higher precision but lower recall in MGCA, or vice versa in MFCL), AMPEND-LS maintains consistently aligned performance across all evaluation metrics. This balanced behavior can be attributed to the complementary integration of agentic multi-persona reasoning, persuasion-aware refinement, and credibility-guided multimodal evidence fusion. While certain baselines such as BREAK perform competitively in specific domains (e.g., MMCoVaR), their performance does not generalize consistently across datasets. In contrast, AMPEND-LS demonstrates stable and uniform gains across heterogeneous domains, suggesting improved robustness under distributional variation. The close alignment among accuracy, precision, recall, and $F1$ further indicates stable decision boundaries and reduced susceptibility to overfitting. Overall, these results establish AMPEND-LS as a robust, adaptable, and evidence-grounded framework capable of effectively addressing diverse misinformation scenarios with enhanced transparency and reliability.
TABLE I: Comparison with SOTA methods
Method AccAcc $F1$ PrePre RecRec PolitiFact LMDD [13] 80.86 75.99 80.05 72.34 MFCL [2] 81.82 82.52 81.61 83.47 BREAK [36] 82.14 82.03 81.79 82.29 MGCA [8] 82.14 81.10 82.14 80.09 EARAM [40] 82.93 81.11 80.90 81.32 FakingRecipe [1] 83.33 80.07 80.83 79.44 FactAgent[16] 88.12 88.51 87.52 88.11 AMPEND-LS (Our) 92.18 ± 1.93 91.64 ± 2.12 92.06 ± 1.85 91.40 ± 2.43 Improvement (%) 4.61 3.53 5.19 3.73 GossipCop LMDD [13] 65.17 71.05 78.21 65.10 FakingRecipe [1] 80.26 80.27 80.21 80.26 MFCL [2] 81.81 81.35 81.65 81.06 BREAK [36] 82.00 82.11 82.51 81.73 FactAgent [16] 83.00 83.00 83.04 83.02 EARAM [40] 85.00 85.49 86.00 85.00 MGCA [8] 88.00 84.09 88.31 80.26 AMPEND-LS (Our) 93.43 ± 1.53 93.24 ± 1.59 93.44 ± 1.48 93.29 ± 1.55 Improvement (%) 6.17 9.07 5.81 9.75 MMCoVaR LMDD [13] 64.62 63.54 78.61 53.40 FakingRecipe [1] 79.96 78.61 79.35 77.92 MFCL [2] 82.06 78.19 77.82 78.57 FactAgent[16] 86.97 85.97 86.33 86.97 MGCA [8] 90.00 88.34 88.19 88.48 EARAM [40] 90.36 88.51 88.32 88.70 BREAK [36] 91.89 89.27 88.73 89.83 AMPEND-LS (Our) 96.46 ± 1.11 95.56 ± 1.31 96.30 ± 1.48 94.96 ± 1.55 Improvement (%) 4.97 7.05 8.53 5.71 Improvement (%) of AMPEND-LS over the second-best method
V-C Ablation Study
To evaluate the contribution of each component in AMPEND-LS, we conduct a systematic ablation study on PolitiFact, GossipCop, and MMCoVaR. The results are summarized in Table II. The key observations are as follows:
(i) Impact of Persuasion Analysis (PKM): The PKM module alone yields relatively modest performance (e.g., $Acc:$ 73.13 on PolitiFact, 67.90 on GossipCop, and 71.94 on MMCoVaR), indicating that persuasion signals by themselves are insufficient for reliable misinformation classification. However, removing PKM from the full pipeline (w/o PKM) results in consistent performance degradation compared to AMPEND-LS (e.g., –6.41 points on PolitiFact, –7.15 on GossipCop, and –10.49 on MMCoVaR in accuracy). This demonstrates that PKM plays a complementary role in resolving uncertain or rhetorically manipulative cases, strengthening final decisions when integrated with evidence and reasoning modules.
(ii) Impact of the Evidence Module: The Evidence-only configuration establishes a meaningful baseline (e.g., $Acc:$ 74.26 on PolitiFact, 79.67 on GossipCop, and 80.04 on MMCoVaR), confirming that credibility-aware retrieval provides substantial factual grounding. However, it remains significantly below the integrated AMPEND and AMPEND-LS models. For instance, on PolitiFact, $F1$ improves from 73.16 (Evidence) to 88.24 (AMPEND) and further to 91.64 (AMPEND-LS). This highlights that reliability scoring, temporal filtering, and multimodal grounding are necessary but not sufficient without structured reasoning and refinement.
(iii) Impact of Agentic Reasoning: The standalone Agentic module outperforms both PKM and Evidence alone (e.g., $Acc:$ 79.92 on PolitiFact, 83.14 on GossipCop, 81.20 on MMCoVaR), demonstrating that structured multi-step reasoning significantly enhances claim analysis. Nevertheless, it still underperforms compared to AMPEND and AMPEND-LS, indicating that reasoning benefits substantially from explicit credibility signals and persuasion-aware refinement. Thus, reasoning alone improves structure, but integration ensures robustness.
(iv) Single-Persona vs. Multi-Persona (ASPEND vs. AMPEND): The single-persona model (ASPEND) achieves strong performance (e.g., $Acc:$ 84.13 on PolitiFact, 88.80 on GossipCop, 86.28 on MMCoVaR), yet consistently trails the multi-persona AMPEND (88.54, 92.68, and 91.27 respectively). This confirms that modeling complementary social roles (journalistic, legal, scientific, supervisory) enhances perspective diversity, reduces bias, and improves decision reliability.
(v) AMPEND vs. AMPEND-LS (Effect of SLM Distillation): Integrating SLM distillation yields further consistent gains across all datasets. For example, on MMCoVaR, accuracy increases from 91.27 (AMPEND) to 96.46 (AMPEND-LS), with corresponding improvements in $F1$, precision, and recall. Similar gains are observed on PolitiFact and GossipCop. These results indicate that the SLM serves as a stabilizing decision layer, mitigating LLM uncertainty while preserving evidence-grounded reasoning. The synergy between agentic reasoning and lightweight classification enhances both performance and deployment scalability.
Overall, the ablation study confirms that each component (credibility-aware evidence retrieval, multi-persona reasoning, persuasion-aware refinement, and SLM distillation) contributes complementary benefits. Their integration yields a system that is not only more accurate but also more stable and robust across heterogeneous misinformation domains.
TABLE II: Module-wise ablation study